### v1

# Fixing non-blocking data structures for better compatibility with memory reclamation schemes

Abstract
We present a new technique, Safe Concurrent Optimistic Traversals (SCOT), to address a well-known problem with Hazard Pointers (HP) related to optimistic traversals. Unlike Epoch-Based Reclamation (EBR), HP enables strong robustness properties but lacks support for some data structures such as Harris’ original list and Natarajan-Mittal tree, which can slow down operations substantially (e.g., when replacing with Harris-Michael list) or make correct implementations impossible altogether (e.g., Natarajan-Mittal tree).
Recent works propose replacing HP with another robust memory reclamation schemes that support optimistic traversals such as HP++. However, HP++ is generally slower than HP, has undesirable applicability trade-offs, and more complex implementation, among other issues. We propose a different method which keeps HP intact but instead relies on data structure adaptations. Moreover, our method applies to other recent (and faster) robust reclamation schemes such as Hazard Eras (HE) and Interval-Based Reclamation (IBR), which are HP-like and have similar problems related to optimistic traversals.
We implement two fundamentally different data structures: Harris’ list and Natarajan-Mittal tree. Our method enables their first correct implementations with optimistic traversals with HP and IBR. Our method is also directly applicable to other data structures which extend Harris’ list, e.g., skip lists and hash maps. We observe similar performance benefits of Harris’ original list (vs. Harris-Michael list) that were previously only available to EBR users. As we discuss in the paper, our Natarajan-Mittal tree is the first bug-free implementation for HP and IBR; it achieves very high throughput, comparable to that of EBR.
Keywords: hazard pointers, non-blocking, Harris’ list, Natarajan-Mittal tree
1 Introduction
With the growing prevalence of multicore and manycore systems, parallelization is becoming crucial for enhancing performance across the entire software stack – applications, middleware, core system services, and low-level components inside an operating system (OS). Non-blocking data structures, which are not dependent on mutual exclusion, are becoming increasingly popular due to their scalability, throughput, and latency characteristics. Unfortunately, many data structures that utilize non-blocking techniques conflict with fundamental assumptions of classical memory management. This is because memory cannot be promptly reclaimed after an object is removed from a data structure, as there may still exist stale pointers to the object due to ongoing operations.
When using manual memory management, memory has to be safely reclaimed by freeing memory only when all ongoing operations with stale pointers complete. A number of past researchers proposed various safe memory reclamation (SMR) techniques in C, C++, or Rust. Non-blocking SMR progress is often crucial for the overall progress guarantees, as we do not want to compromise data structure’s progress properties. Unfortunately, many SMR schemes are trying to attain several highly desirable properties simultaneously while often unknowingly compromising non-blocking progress or other important properties one way or the other. A recent work [21] sheds more light onto this conundrum: it proves that it is impossible to achieve three highly desirable SMR properties simultaneously, i.e., only two properties can ever be achieved: (A) robustness (bounded memory usage), (B) easy integration (no dependency on any special APIs, such as roll-back mechanisms or OS-based primitives), and (C) wide applicability (compatibility with most non-blocking data structures).
While (A) is non-negotiable for non-blocking progress, [^1] If memory usage is unbounded, no further system progress can be made once memory is exhausted, regardless of whether actual locks are present. past researchers would compromise either (B), (C), or other properties that are tangentially related to non-blocking progress, sometimes without fully realizing that. To make matters worse, many researchers prioritize (C) because, intuitively, their goal is to support as many data structures as possible. However, we argue that sacrificing (C) rather than (B) by carefully adapting existing problematic data structures is a “necessary evil”, which is a crucial new insight. First, (C) in its strongest form – applying to all data structures – only currently known to exist in blocking SMR schemes [21], which is a non-starter for us. Second, even if we accept (C) in its relaxed form – applying to most data structures – the absence of (B) would lead to either inability to achieve non-blocking progress in practice due to dependence on OS-based primitives, lack of universally-defined APIs, or both. (We discuss some specific examples in Section 5.) Third, the absence of (C) simply implies that adaptations are required, not impossibility of the implementation.
EBR [9] is a fast and easy-to-use approach but has a serious drawback that any stalled thread causes unbounded memory usage, i.e., it lacks (A). Once exhausting memory, the system must either reboot or block until memory is freed. Many researchers prefer EBR’s simplicity but attempt to solve its robustness issue. Unfortunately, such approaches typically sacrifice strict non-blocking progress [5, 22, 4, 1, 2], as discussed in Section 5, or need non-trivial roll-back mechanisms [20], which are fairly hard to generalize in terms of the API within the same library framework.
HP [14] is an antipode of EBR: it is somewhat slower and more difficult to use but provides robustness with strict non-blocking guarantees by placing a hard limit on the number of local pointers within each thread. Prior to this work, well-known data structures such as Harris’ linked list [10] with optimistic traversals [^2] The original paper [10] did not explicitly mention that search can be done in a read-only manner (w/o any CAS operations). In this paper, by Harris’ list, we denote the enhanced version [11] with optimistic read-only traversals. and Natarajan-Mittal’s binary search tree (BST) [15] were not known to work [^3] Despite presenting Natarajan-Mittal Tree results in [25, 17] for HP and other SMR schemes, it was later argued [3] that such implementations are buggy and leak memory. Moreover, [3] pointed out two bugs, and only one bug was resolved in [3]. Prior to our work, the second bug related to optimistic traversals remained unresolved. with HP [12, 3], which is due to HP’s lack of the wide-applicability (C) property. A number of other recent schemes, including Interval Based Reclamation (IBR) and Hazard Eras (HE), sometimes offer better performance but have similar issues to HP. HP++ [12] is another recent scheme. HP++ is slower than HP and, strictly speaking, makes compromises with both (B) and (C). However, HP++ still demonstrates support for these two data structures. We are inspired by HP++’s success with these data structures, which is indicative that many existing data structures can still be implemented, with an additional effort, despite prior beliefs regarding limited applicability of the SMR schemes in this category.
In this paper, we depart from a typical strategy of designing a “silver-bullet” SMR approach for the reasons described above. Instead, we argue that the existing problematic data structures can be modified to accommodate SMR schemes with strict non-blocking properties. We demonstrate our new method in Section 4 for Harris’ list, Natarajan-Mittal tree, and other data structures. In Section 5, we also touch upon how limitations of our approach with respect to wait-free traversals can be addressed.
2 Background
2.1 Lock- and Wait-Freedom
The most widely used non-blocking progress is lock-freedom, which allows threads to interfere with each other while still guaranteeing that at least one thread makes progress in a finite number of steps. This should not be confused with “lockless” approaches that simply avoid explicit locks, where one preempted thread can block all other threads. Wait-freedom is a much stronger property which requires that all threads make progress in a finite number of steps.
2.2 Memory Reclamation API
For both EBR and HP, the API is called directly. Each thread maintains a reservation, which is a globally accessible per-thread state. Memory objects, containing the required SMR header, are allocated using standard OS methods. Once a memory object is part of a data structure, it can only be reclaimed through two steps. First, after removing a pointer from the data structure, a memory object is retired. When the object becomes unreachable by any other concurrent thread, it is then returned to the OS.
EBR requires specifying both the start and end points of an operation to define the scope of protected memory accesses. In contrast, HP offers finer granularity in terms of reserved objects, requiring to explicitly wrap accesses to shared-memory objects. Additionally, in HP, the number of local pointers that can be held simultaneously by each thread is limited, and local pointers must be differentiated by unique indices.
2.3 Overview of Data Structures
In Table 1, we characterize common data structures based on their speed, (original) compatibility with SMR schemes, and compatibility when using our method. We omit hash maps as they are mostly arrays of Harris’ or Harris-Michael lists.
Table 1: Compatibility and performance characteristics of common non-blocking data structures.
Data Structure Performance EBR Original HP/IBR/HE HP/IBR/HE with SCOT
Harris’ Linked List [10]
Harris-Michael Linked List [13]
Fraser’s Skip List [9]
Herlihy-Shavit Skip List [11]
Natarajan-Mittal Tree [15]
Ellen et al. Tree [8]
These data structures employ “logical” deletion, i.e., a node inside a data structures is first simply marked (by stealing one bit from a node pointer), and then unlinked from the data structures (i.e., “physically” deleted). The original Harris’ List and Fraser’s Skip List use optimistic traversals, i.e., a chain of consecutive logically deleted nodes is deleted using one CAS operation; they allow implementing read-only search operations, which simply skip logically deleted nodes when iterating through the data structure. These optimizations improve performance and reduce contention.
However, as we discuss below, this approach is incompatible with HP. Harris-Michael List and Herlihy-Shavit Skip Lists use a different approach: they still use “logical” deletion. However, they mandate that logically deleted nodes are removed immediately after the first encounter and before proceeding further, including the search operation.
Ellen et al. Tree and Natarajan-Mittal Tree are more complex algorithms and typically achieve a higher overall throughput. There is a similar difference between them with respect to optimistic traversals. However, Natarajan-Mittal Tree’s ability to remove multiple logically deleted tree branches at the same time is even more crucial here – there is a higher chance of contention when multiple threads access the same node due to the hierarchical (rather than linear) structure of trees.
As we show in the table, data structures with optimistic traversals are incompatible with HP, IBR, and HE, a problem that we address in this paper.
2.4 Harris’ and Harris-Michael Linked Lists
The main challenge in lock-free linked lists is allowing a deleting thread to mark a node as deleted and update the preceding node’s next pointer. Timothy L. Harris solved this [10] by using helper threads during deletion. A deleting thread first marks a node as “logically” deleted by updating its next pointer, and then the node is “physically” removed by the deleting thread or a helper (conflicting concurrent operation). Harris’ list supports optimistic traversals, meaning that logically deleted nodes can still be traversed and bypassed by a search operation without immediate physical removal. This allows threads to defer unlinking logically deleted nodes until reaching the intended destination node.
However, this approach presents fundamental challenges for HP. Figure 1 shows an example with several concurrent operations. Suppose Thread 3 (not shown in the figure) has previously marked N2 for deletion but has not yet unlinked it. Then, Thread 2 marks N3 for deletion and unlinks the entire chain of nodes between N1 and N4 (assuming all consecutive nodes in the chain are logically deleted). Meanwhile, Thread 1 is traversing the list; it reaches N2 before Thread 2 physically removes the chain. Although N2 is now retired by Thread 2, it still remains in the HP limbo list. The reason for that is that Thread 1 has made a reservation prior to its physical deletion. N3 is also retired by Thread 2, but because Thread 1 never reserves N3, it is not guaranteed to stay in the HP limbo list, i.e., N3 is returned to the OS. Consequently, Thread 1 fails when accessing N3.
Maged M. Michael modified [13] Harris’ approach to resolve compatibility issues with HP: when a thread encounters a node marked as logically deleted during traversal, it immediately attempts to physically remove the node from the list.
Why does Michael’s approach work? Recall that accessing N2 is still safe, because its HP reservation is made prior to N2’s physical deletion. (To successfully make a reservation, HP verifies that the pointer from N1 to N2 remains intact.) Once N2 is physically deleted, it is retired and ends up in the HP limbo list. If we now attempt to reserve N3, HP falsely succeeds as the pointer from N2 to N3 remains intact unlike the pointer from N1 to N2. Thus, Michael’s approach is to ensure that pointers always change during physical removal, which happens automatically when we remove one node at a time. [^4] Another way to think about Michael’s approach is that if we remove nodes one by one, N3 takes place of N2, i.e., right after N1. N3 will be the first logically deleted node in the chain when making the HP reservation, which is still safe to access.
While it fully solves the HP problem, Michael’s approach increases the number of CAS operations and may lead to a higher contention among threads. It also makes read-only traversals impossible.
Figure 1: Unsafe optimistic traversal of Harris’ linked list with HP: accessing the first logically deleted node (N2) is safe but subsequent nodes may cause SEGFAULTs.
1 struct {
    node_t * Next;
     // Next node pointer
    key_t Key;
     // Any key type
2
3    } node_t;
   node_t * Head;
     // Head of the list
4    void Init()
5        node=malloc(sizeof(node_t));
6        node->Next = nullptr;
7        Head = node;
8
9   bool Insert(key_t key)
10        new = malloc(sizeof(node_t));
11        new->Key = key;
12        node_t **prev, *curr, *next;
13        while true do
14            if ( Do_Find(key, &prev, &curr, &next, false) )
15                free(new);
16                return false;
17
18           new->Next = curr;
19            if ( CAS(prev, curr, new )  return true;
20
21
22   bool Delete(key_t key)
23        node_t *curr, *next;
24        while true do
25            if ( !Do_Find(key, &prev, &curr, &next, false) )
26                return false;
27
28           if ( !CAS(&curr->next, next, getMarked(next) )  continue;
29            if ( CAS(prev, curr, next )  smr_retire(curr);
30            return true;
31
32
28 bool Do_Find(key_t key, node_t *** p_prev, node_t ** p_curr, node_t ** p_next, bool is_search)
29    node_t **prev, *curr, *next;
30    prev_next = nullptr;
31    prev = &Head;
32    curr = Head;
33    while true do
34        if ( curr == nullptr )  break ;
35        next = curr->Next;
36        if ( !isMarked(next) )
37            if ( curr->Key $\geq$ key )  break ;
38            prev = &curr->Next;
39            prev_next = next;
40
41       curr = getUnmarked(next);
42
43   if ( !is_search && prev_next != curr ) // Skip for Search
44        if ( !CAS(&prev, prev_next, curr )  goto 2 ;
45        do // Retire nodes from the deleted chain
46            node_t *node = getUnmarked(prev_next->next);
47            smr_retire(prev_next);
48            prev_next = node;
49
50       while (prev_next != curr);
51   *p_curr = curr;
52    *p_prev = prev;
53    *p_next = next;
54    return curr && (curr->Key == key);
55
56bool Search(key_t key)
57    node_t **prev, *curr, *next;
58    return Do_Find(key, &prev, &curr, &next, true);
59
Figure 2: Harris’ Linked List with optimistic traversals (ignoring memory reclamation).
In Figure 2, we demonstrate Harris’ list without the HP reclamation. The Insert method inserts a new node into the list. Before inserting a new node, Insert checks if the key of the to-be-inserted node exists via Do_Find. The new node gets inserted unless its key is already present in the list. Delete removes a node from the list. If the key is found in the list, the node is logically deleted at L25 by marking next, and then one attempt is made to unlink it from the list at L26.
In Search, the underlying Do_Find operation skips L41. In Insert and Delete, we also need to locate the right position for the key. Moreover, Insert and Delete ensure that we are not modifying pointers nodes which belong to nodes which are already (logically) deleted by cleaning up the discovered chain of logically deleted nodes in L41-L47 first. If we simply integrate HP without any modifications, L44 may crash while retrieving the next pointer in a loop.
3 Natarajan-Mittal Tree
Natarajan-Mittal Tree [15] is a fairly complex algorithm and its detailed discussion and pseudocode is beyond the scope of this paper; we describe its high-level idea, which suffices to explain how to apply SCOT to this tree. The tree keeps left and right pointers in every node. The actual keys and values are stored in leaf nodes, while internal nodes are simply guiding in which direction (left or right) the underlying search operation has to descend to the next level. A concept similar to logical deletion also exists in this tree. There are two types of marking: “flagging” when marking the leaf node for its logical deletion followed by “tagging” of its sibling node.
When performing the underlying search operation, we keep the successor node which is the last untagged node as well as its preceding node – ancestor. We also keep the immediate parent of the leaf node. The algorithm makes a very crucial observation: a chain of consecutively tagged edges can be eliminated with one CAS operation by updating ancestor’s link from successor to the remaining leaf node as shown in Figure 3.
Figure 3: Natarajan-Mittal tree: entire tagged edges can be removed.
4 Design
4.1 Bird’s-Eye View
The crux of the problem with the HP implementation is its inability to properly track physically deleted nodes in the original Harris’ list (or tagged edges in Natarajan-Mittal tree) while traversing through logically deleted nodes. However, if there was some way to confirm that the following logically deleted node is not yet physically deleted, optimistic traversals are still feasible.
In Figure 1, N2 is still safe to access because it is not physically deleted, i.e., N1 still points to N2. Moreover, according to Figure 2, we always make sure that the node to the left is not logically deleted and thus never unlink nodes in the middle of the chain, i.e., we can delete just N2, or both N2-N3, but we never delete N3 while keeping N2 in the list. In other words, we either remove the entire chain of logically deleted nodes or a subchain with the truncated tail.
Combining these observations together, we make one crucial insight that helps solve the issue with HP: It is still safe to access N3 and the following nodes as long as at every step, i.e., after making an HP reservation for the next logically deleted node, we verify that N1 still points to N2. In other words, once we reach N2, we declare it to be a “dangerous zone”. Until we reach the end of the chain of consecutive logically deleted nodes, we are going to stay in the dangerous zone and perform additional checks. More specifically, after retrieving N3’s pointer and creating its HP reservation, we check that N1 still points to N2. We will continue to perform these checks at every iteration until we reach the end of the chain. If the check fails, we cannot proceed further and need to restart the search operation from the very beginning.
4.2 SCOT for Harris’ List
In Figure 4, we present our approach for Harris’ list. Compared to Michael’s approach, we need an additional hazard pointer index (Hp3) which protects the first unsafe, i.e., logically deleted, node. This extra hazard pointer prevents the ABA problem, i.e., a case when the unsafe node gets recycled while we are traversing the list. By holding this extra hazard pointer, we can solely rely on pointer comparison. In other words, Hp3 protects N2 from our example above irrespective of where we currently are in the chain of logically deleted nodes. Like in Michael’s approach, we need Hp2, which protects the prev area. We note that in Harris’ list, it will be the last safe, i.e., not logically deleted, node.
The HP reclamation provides protect to safely retrieve the object and make a corresponding reservation. Additionally, we implement duplicate to duplicate the existing hazard pointer when moving to the next iteration. It is crucial to duplicate hazard pointers such that the old HP index has a lower numerical value than the new HP index (e.g., Hp0 to Hp1). This allows to avoid a small race window when iterating in the same (ascending) order of indices in retire.
Once we get to the dangerous zone of logically deleted nodes, Figure 4 runs a check in L16 which will ensure that reservation made in L8 is correct. Once the chain (or its subchain) gets unlinked, L16 fails and we restart the operation from the very beginning. We highlight in yellow color all major changes related to SCOT in HP.
1 bool Do_Find(key_t key, node_t *** p_prev, node_t ** p_curr, node_t ** p_next, bool is_search)
2    node_t **prev, *curr, *next;
3    prev_next = nullptr;
4    prev = &Head;
5    curr = hp.protect(Head, Hp1);
6    while true do
7        if ( curr == nullptr )  break ;
8        next = hp.protect(curr->Next, Hp0);
9        if ( !isMarked(next) )
10            if ( curr->Key $\geq$ key )  break ;
11            prev = &curr->Next;
12            prev_next = next;
            hp.duplicate(Hp1, Hp2);
             // Duplicate Hp1 (curr) to Hp2
              hp.duplicate(Hp0, Hp3);
             // Duplicate Hp0 (next) to Hp3
13
14       else
            // Dangerous zone (marked): Validate that last safe node (N1) still points to the first unsafe (N2) node
15              if ( *prev != prev_next ) goto 4;
16
17       curr = getUnmarked(next);
        hp.duplicate(Hp0, Hp1);
         // Duplicate Hp0 (next) to Hp1
18
19
// <—— Continuation of Do_Find() ——>
19
20    if ( !is_search && prev_next != curr ) // Skip for Search
21        if ( !CAS(&prev, prev_next, curr )  goto …;
22        ;
23        do // Retire nodes from the deleted chain
24            node_t *node = getUnmarked(prev_next->next);
25            smr_retire(prev_next);
26            prev_next = node;
27
28       while (prev_next != curr);
29   *p_curr = curr;
30    *p_prev = prev;
31    *p_next = next;
32    return curr && (curr->Key == key);
33
34
// Hazard Pointer Indices
const int Hp0 = 0;
  // Protects next
const int Hp1 = 1;
  // Protects curr
const int Hp2 = 2;
  // Protects last safe node (prev)
  const int Hp3 = 3;
  // Protects first unsafe node
Figure 4: SCOT for Harris’ list using HP reclamation: showing a new version of Do_Find only.
Figure 5: SCOT for Harris’ list: validating *prev = prev_next at every iteration while traversing through the dangerous zone.
4.3 SCOT for Natarajan-Mittal Tree
For the tree, we allocate 5 hazard pointers to protect nodes in the underlying search procedure: current, leaf, parent, successor, and ancestor. The current node points to the lowest node that is currently considered.
To properly support optimistic traversals, after each HP reservation of the current node, if the corresponding node is flagged or tagged, we need verify that ancestor still points to successor. If ancestor points to some other node or successor becomes tagged, we need to restart from the very beginning.
This idea is largely similar to the above-mentioned Harris’ list approach, except that no additional hazard pointer is required, as we already need to protect both ancestor and successor nodes. Additionally, we need to check both tagged and flagged conditions for the current node, as (flagged) leaf nodes’s hazard pointer reservations need to be also verified.
Due to complexity of the tree algorithm and very similar checks for optimistic traversals, we omit full pseudocode changes in this paper.
4.4 SCOT for Hash Map and Skip Lists
SCOT for these data structures is largely identical to Section 4.2. Hash Maps are directly based on linked lists. Fraser’s skip list uses the same idea of logically deleted nodes as Harris’ list.
5 Related Work
5.1 Non-Blocking SMR
The ERA theorem [21] demonstrates that at most two out of three properties – (A) robustness, (B) easy integration, and (C) wide applicability – can be achieved simultaneously. Moreover, strong applicability, as defined by that work, currently seems to be available only with SMR schemes that implement some kind of quiescence, similar to EBR. However, quiescence periods inevitably lead to potentially unbounded memory usage (i.e., lack of non-blocking progress).
A number of schemes, e.g., DEBRA+ [5], QSense [4], ThreadScan [2], ForkScan [1], NBR [22], propose a solution to EBR’s robustness (A) problem. However, they all need special OS-based support such as interaction with the OS scheduler or signal delivery. These OS mechanisms inevitably use locks in most typical OSs. They satisfy (A) and (C) properties but not (B), which is essential to the strict non-blocking progress. Other schemes that satisfy (A) and (C), e.g., VBR [20], need non-trivial roll-back mechanisms, which are too complicated to generalize in terms of the uniform library API.
There were also recent attempts to build non-blocking data structures using automatic garbage collection techniques. Generally speaking, garbage collection is subject to the same trade-offs and limitations, discussed above. FreeAccess [6] addresses the problem with (C) and optimistic traversals but forgoes (B) by requiring to divide code into separate read and write phases. OrcGC [7] is another lock-free garbage collector. OrcGC only works with data structures without loops and requires mechanisms similar to smart pointers, either in the compiler or the language itself; it can also be slower than HP in certain tests.
Other SMR schemes such as interval-based reclamation (IBR) [25], hazard eras (HE) [19], wait-free eras (WFE) [16], Hyaline [17], and Crystalline [18] provide HP-like interface; they still lack the wide-applicability property (C) but are often faster than HP. The method proposed in this paper to mitigate the problem with (C) is suitable to all of them.
5.2 Non-Blocking Data Structures
Timothy L. Harris [10] was the first to present a practical non-blocking lock-free linked list. Maurice Herlihy and Nir Shavit pointed out [11] that read-only optimistic search operations are feasible with Harris’ list. Unfortunately, Harris’ list is incompatible with HP [7, 6, 12, 3, 18, 17]. Maged M. Michael proposed [13] to disable optimistic traversals by unlinking logically deleted nodes one by one, making it feasible to use the list with HP. Our work allows to support Harris’ original list with HP and take advantage of optimistic traversals.
A skip list is a data structure with an efficient search, insertion, and deletion operations, achieving (probabilistically) logarithmic time complexity. Lock-free versions of skip lists [11, 9] are very similar to Harris’ list and Harris-Michael list, except that they effectively consist of several sublists. Consequently, challenges related to logical deletion and optimistic traversals are identical.
Hash maps store and retrieve key-value pairs in constant time. A lock-free hash map [13] is, effectively, an array of linked lists, which can be based on either Harris’ or Harris-Michael approach.
Binary search trees are also commonly used for efficient search, insertion, and deletion operations. Ellen at al. Tree [8] implements one such approach which is compatible with HP. However, Natarajan-Mittal Tree [15] is significantly faster (around 43% according to [12]) due to its ability to remove multiple concurrently tagged edges in one pass. Natarajan-Mittal Tree is a very popular data structure and evaluated in several recent works [25, 3, 17, 12]. Prior to our work, as pointed out in [3, 12], Natarajan-Mittal Tree was incompatible with HP, IBR, and other similar schemes.
5.3 Wait-Free Optimistic Traversals
With our approach, optimistic traversals might have to occasionally restart the search operation from the beginning if, due to overlapping modifications, the state of the data structure diverges significantly and cannot be recovered locally. This guarantees lock-free but not wait-free progress. This limitation is very similar to that of HP++’s existing solution which also does not support wait-free optimistic traversals due to potential restarts.
However, this restart is inherently unavoidable with data structures that use optimistic traversals: it is still needed for insertion and deletion operations, regardless of the search operation or the reclamation scheme used. We note that we can still implement wait-free search traversals by using Timnat-Petrank’s method [23, 24] which allows restarts and works with HP. This method can be applied to various data structures, including linked lists. Timnat-Petrank’s original wait-free linked list, implemented in HP, is similar to Harris-Michael list. With our method, it can now leverage Harris’ optimistic traversals for faster performance.
6 Conclusion
We introduced safe concurrent optimistic traversals (SCOT), a new method which enables support for data structures that are normally incompatible with HP, IBR, HE, and other similar methods that lack the wide applicability support. Despite prior beliefs [12] of incompatibility of Natarajan-Mittal Tree and Harris’ List with HP, we have found that not only they are feasible with HP, IBR, and HE, but they also can largely retain performance benefits when comparing to equivalent data structures without optimistic traversals.
We hope that our work will help to spur further research and reevaluate the existing challenges related to safe memory reclamation because lack of the wide applicability property may be addressed by other means.
Availability
SCOT’s code and benchmark for Harris’ list and Natarajan-Mittal tree is available at https://github.com/rusnikola/scot